High school seniors with strong academic records apply to the nation's most selective colleges in greater numbers each year. Because the number of slots remains relatively stable, some colleges reject more early applicants. Suppose that for a recent admissions class, an Ivy League college received 2,851 applications for early admission. Of this group, it admitted 1,030 students early, rejected 857 outright, and deferred 964 to the regular admission pool for further consideration. In the past, this school has admitted 18% of the deferred early admission applicants during the regular admission process. Counting the students admitted early and the students admitted during the regular admission process, the total class size was 2,372. Let E, R, and D represent the events that a student who applies for early admission is admitted early, rejected outright, or deferred to the regular admissions pool.

(a)

Use the data to estimate

P(E), P(R), and P(D).

(Round your answers to four decimal places.)

P(E

P(R)

P(D)

(b)

Are events E and D mutually exclusive?

They  ---Select--- are or are ,not mutually exclusive.

Find

P(ED).

P(ED)

=

(c)

For the 2,372 students who were admitted, what is the probability that a randomly selected student was accepted during early admission? (Round your answer to four decimal places.)

(d)

Suppose a student applies for early admission. What is the probability that the student will be admitted for early admission or be deferred and later admitted during the regular admission process? (Round your answer to four decimal places.)
Answer #1

a)

P(E)=1030/2851= 0.3613
P(R)=857/2851= 0.3006
P(D)=964/2851= 0.3381

b)

Yes they are mutually exclusive
P(E n D) =0

c)

c)probability =1030/2372= 0.4342

d)

probability =(1030+0.18*964)/2851= 0.4221